OpenStudy (anonymous):

Prove whether it's true or find a counterexample to demonstrate that it is false: For any function f, f(x+y)=f(x)+f(y).

OpenStudy (anonymous):

it is obviously not true in general. for example, \[\sin(x+y) \neq \sin(x) + \sin(y)\]

OpenStudy (anonymous):

alright, thanks.